March 12, 2011

LUCIFER 1 Mask exchanger failed yesterday morning. Germans worked on it all day, but do not know what is wrong. We decided to do only imaging tonight.

Big earthquake today in Japan.

Observers: Wiphu Rujopakarn, Jill Bechtold, Dave Thompson

cirrusy at dusk

Primary mirror 1.3 deg warmer than ambient at 18 deg. twilight.

02:00UT finished taking flats (K x 3, Ks x 3, J, H; img # 2-103). Some clouds in twilight sky.

02:02UT executed MS0451 imaging script in K. Some passing cirrus. Seeing ~1".

02:47UT executed MS0451 imaging script in H.

03:19UT re-executed MS0451 imaging script in K. Frequent passing cirrus. Seeing > 1".

03:25UT lost guide star due to cloud.

03:34UT re-execute imaging script in K (starting img#132). Frequently lose guide star due to clouds.

03:53UT execute 3C220 imaging script in K. Frequent passing cirrus.

05:10UT execute 3C220 imaging script in H (starting img#167)

05:48UT execute 3C220 imaging script in J (starting img#182). Thick cirrus.

05:57UT lost guide star due to cloud; lost pressure on primary mirror. Stopped script.

06:08UT re-execute 3C220 imaging script in J (starting img#185)

06:54UT execute MS1149 imaging script in K, and subsequently J. Seeing ~0.8" - >1" in guide camera

08:52UT execute MS1149 imaging script in H (starting img# 245). Seeing ~0.6" in guide camera

09:37UT David is worried about wind, and is moving the dome around

09:39UT execute sdss1343.K.txt. (starting img #260) 15x120s. Started first exposure at UT 9:43. seeing in guide camera .55

clouds

10:07UT seeing blowing up to 2" currently reading out #269. rats.

10:18UT seeing back down to .5" in guide camera.

10:21UT (3:20am) Started sdss1343.H.txt. 10x120s.

10:48UT (3:48am) Started sdss1343.J.txt 10x120s. Seeing in Guide Camera .48"

11:21UT execute imaging script for A2218 in J. ~1.0" seeing in guide camera (starting img# 295). Relatively clear sky. 0.7"-0.9" seeing.

12:06UT abort J-band script; the wind is getting strong (16.4 m/s). David rotates the dome around to do windscanning.

12:12UT ====== Astronomical Twilight =========

12:18UT windscanning completed Back on target. Max wind reaches 17.1 m/s. execute imaging script for A2218 in K. Guider seeing 0.8" - 1.2"

12:31UT sky brightness in R band reaches 21.0 mag/arcsec^2

12:44UT wind reaching 18.4 m/s. Close dome. R sky brightness ~ 19.5 mag/arcsec^2

Dark frames starting img#322
Topic revision: r2 - 12 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback